HIGH RESOLUTION SPECTROSCOPY OF HEXAMETHYLENETETRAMINE (HMT) C_{6}N_{4}H_{12}

\underline{V. BOUDON}, Laboratoire Interdisciplinaire Carnot de Bourgogne, UMR 6303 CNRS-Université de Bourgogne, 9. Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; O. PIRALI^{a}, Ligne AILES

– Synchrotron SOLEIL, L’Orme des Merisiers, F-91192 Gif-sur-Yvette, France.

Hexamethylenetetramine, or HMT (C_{6}N_{4}H_{12}) is a N-substituted derivative of adamantane C_{10}H_{16} which is the smallest sample of the diamondoid molecules family. Thanks to their high stability, diamond-like molecules have long been suspected to be present in space^{b} (note that diamond nanocrystals are extracted from Murchinson meteorites^{c}), and HMT is known to be an abundant residue of UV irradiated ice analogs^{d} and might be present in Titan’s atmosphere. Using the Bruker IFS 125 coupled to a multipass cell (absorption path length of 150 m) of the AILES beamline at SOLEIL, we recorded the IR spectrum of gas phase HMT in the 300–3000 cm^{−1} spectral region with an unapodized resolution 0.001 cm^{−1}. HMT is a solid powder with about 0.008 mbar vapour pressure at room temperature, it is a T_{d} molecule (as adamantane) and has 25 vibrational modes from which only 9 are infrared active. Over the 9 IR active modes, we were able to rotationlly resolved the spectra of 6 of them. The analysis of all the resolved bands has been performed thanks to the XTDS and SPVIEW softwares developed in Dijon for such molecules^{e}. Each band can be considered as isolated and we get very good fits of line positions, with a root mean square deviation better than 5 × 10^{−4} cm^{−1} for J values up to 80 or more in each case. As for our recent study concerning adamantane^{f}, the resulting synthetic spectra will permit an active search of this very stable specie in different sources of the interstellar medium.

aAlso at: Institut des Sciences Moléculaires d’Orsay, UMR8214 CNRS-Université Paris-Sud, Bat.210, 91405 Orsay cedex, France.

bW. C. Saslaw and J. E. Gaustad, Nature, 221, 160 (1969)

cR. S. Lewis et al., Nature, 326, 160 (1987)

dM. P. Bernstein et al., ApJ, 454, 327 (1995)

eCh. Wenger, V. Boudon, M. Rotger, M. Sanzharov and J.-P. Champion, J. Mol. Spectrosc., 251 102–113 (2008).

fO. Pirali, V. Boudon, J. Oomens, M. Vervloet, J. Chem. Phys., 136, 024310 (2012)